A man has to go 50 m due north, 40 m due east and 20 m due south to reach a field. (a) What distance he has to walk to reach the field? (b) What is his displacement from his house to the field?

Grade:10

2 Answers

Jitender Pal
9 years ago
(b) AF = AB – BF = AB – DE = 50 – 20 = 30 m His displacement is AD AD = √(AF^2-DF^2 ) = √(〖30〗^2+〖40〗^2 ) = 50m In ∆AED tan θ = DE/AE = 30/40 = ¾ ⇒ θ = tan-1 (3/4) His displacement from his house to the field is 50 m, tan-1 (3/4) north to east.
Gaurav Kumar
6 years ago
Gjhgufyhduhfhu8tt7665ufyifgjvxyiusjrhrirt jghrurh4ui8rhrvdi3j4bejei2u3ie ki yudhfd du gaurav l kumar krib vyngy